OpenStudy (smurfy14):

log(base5)(x-4)^2-log(base5)3=log(base5)3 help please!!!!!

OpenStudy (anonymous):

\[\log_{5} (x-4)^{2} - \log_{5} 3=\log_{5} 3\] \[2\log_{5} (x-4)= 2\log_{5} 3\] x-4=3 x=7

OpenStudy (smurfy14):

awsome thanks!! any ideas on lnx=3.612? i got 37.040 but idk if thats right?

OpenStudy (anonymous):

looks good! check with calculator value of: e^(3.612)